Given y=1/2x-1, which of the following tables of value are correct?
Table A
x,y
2,0
0,-1
-2,-2
Table B
x,y
5,2
2,0
-5,-2
Table C
x,y
6,3
7,2
-8,-3
Table D
x,y
1,1
-1/2, -1/2
-1,-1

Answer by KnightOwlTutor(293)

Just plug in the x value. The table A is the only one that is the correct solution for y.
y=1/2x-1 0=(1/2)(2)-1